# NSD-DIL: Null-Shot Deblurring Using Deep Identity Learning

### v1

@@ 4.1 Training setup
We trained the proposed L-CNN (refer Sec.3.3 for architecture details of L-CNN) on the RKG dataset (refer Sec. 3.2) with the learning objective given in Eq. (4). The number of epochs was $40$ and the learning rate was $410e-$ 4. The Adam [22] optimization was used, with $\beta=0.9$. The values of hyper-parameters used in Eq. (5), set empirically, are as follows, $\lambda_{1}=0.8,~{}\lambda_{2}=0.8,~{}\lambda_{3}=0.4$. The NSD-DIL code and RKG dataset will be made publicly available soon for reproducibility.
@@ 4.2 Computational complexity:
The proposed L-CNN is an efficient and lightweight deblurring model. The computational complexity in terms of parameters and inference time of the NSD-DIL model was compared to SotA methods and is given in Table 1. Here, the proposed NSD-DIL model requires very less computational resources, by order of 1000 in terms of parameters and inference time when compared to the recent SotA work i.e. IFFS [56]. Besides NSD-DIL, the deblurring task with DRK further reduces the required computational resources parameters and inference time by an order of 10 to NSD-DIL and by an order of 10K to SotA methods. Thus the proposed NSD-DIL model is more suitable for real-time deblurring tasks.
@@ 4.3 Results
The deblurring performance of our proposed method is compared with a wide range of methods including standard and recent non-DL blind deblurring methods NSM [24], PIBKE [36], Polyblur [10] and DL-based methods DeblurGAN [25], DUN [61], XY [18], IFFS [56]. For the deblurring task, the blind deblurring methods, NSM proposed a regularization term in the spatial domain based on normalized sparsity measure. While, the PIBKE and IFFS operate on the frequency domain features of the input blurred image for $K$ estimation based on the phase information of the blurred image and frequency filtering techniques respectively. The methods DeblurGAN and XY are supervised and DUN is an unsupervised data-driven deblurring method. We note that the proposed NSD-DIL is only the DL approach among the above-mentioned DL-based methods that achieve the deblurring task without training on image data.
@@
NSM PIBKE DeblurGAN DUN XY Polyblur IFFS NSD-LCNN NSD-DRK
@@
120 120 1 1 0.729 0.001 3.5 0.0002 0.0005
$K_{11\times 11}$ 27.02/0.8053 24.19/0.7280 26.71/0.7925 24.32/0.7400 27.32/0.8097 27.21/0.7829 26.59/0.8276 27.26/0.8434 28.09/0.8405
$K_{15\times 15}$ 27.63/0.8227 24.83/0.7312 26.65/0.7940 24.32/0.7390 27.12/0.8041 27.06/0.7771 26.50/0.8225 26.96/0.8350 27.99 /0.8384
$K_{19\times 19}$ 27.66/0.8232 24.89/0.7439 26.68/0.7970 24.33/0.7428 27.30/0.8111 27.20/0.7874 26.51/0.8260 26.97/0.8328 27.94/0.8385
$K_{23\times 23}$ 27.57/0.8162 25.44/0.7552 26.59/0.7911 24.31/0.7384 27.07/0.8027 26.96/0.7707 26.51/0.8215 26.95/0.8333 28.05/0.8372
$K_{27\times 27}$ 27.67/0.8176 24.97/0.7341 26.62/0.7921 24.29/0.7366 27.10/0.8019 27.11/0.7789 26.53/0.8235 26.90/0.8315 28.02/0.8373
PSNR: 27.51 $\pm$ 0.2474 24.86 $\pm$ 0.3997 26.65 $\pm$ 0.0424 24.31 $\pm$ 0.0135 27.18 $\pm$ 0.1059 27.10 $\pm$ 0.0928 26.52 $\pm$ 0.0324 27.01 $\pm$ 0.1167 28.02 $\pm$ 0.0511
SSIM: 0.8170 $\pm$ 0.0064 0.7384 $\pm$ 0.0099 0.7933 $\pm$ 0.0020 0.7396 $\pm$ 0.0020 0.8059 $\pm$ 0.0037 0.7794 $\pm$ 0.0056 0.8242 $\pm$ 0.0022 0.8333 $\pm$ 0.0012 0.8383 $\pm$ 0.0011
@@
Comparison on simulated blur: We generated a synthetic mild blur dataset following the [10]. Random Gaussian kernels of varying shapes, orientations ) (i.e. $\sigma_{1},\sigma_{2}\in U[0.175,3]$ and rotation angle $\theta\in U[0,\pi]$) and sizes $11\times 11,15\times 15,19\times 19,23\times 23,27\times 27$ were generated. Each set of kernels, varying in size, was applied to the set of 100 sharp images from the DIV2K validation dataset [52]. To have a little deviation from the Gaussian distribution, a small amount of multiplicative noise was introduced to the random Gaussian kernels. We compared the NSD-DIL model to the above-mentioned SotA works using the simulated blur dataset. To assess the ability to restore the finer details during the deblurring process, we employ PSNR and SSIM for qualitative and quantitative similarity with the corresponding sharp images. The results were tabulated and presented in Table 1. The qualitative results were provided in Fig. 4.
Despite being very lightweight, the proposed NSD-DIL demonstrated its effectiveness significantly with standard metrics. The compared SotA methods need supervised learning from images for $K$ estimation and/or deblurring model learning. Thus requires more memory, and time resources at least by order of 100 in parameters and inference time respectively compared to the proposed NSD-DIL.
Evaluation on real data: Further, we provide the qualitative comparison of deblurring performance on real captured blurred images provided in Fig. 1. The qualitative results given in Fig. 4 and Fig. 1 illustrate that the proposed model did not produce artifacts in the generated deblurred output images, unlike the compeer, zero-shot non-DL SotA works like [24], [36]. Besides, the proposed NSD-DIL method demonstrates its generalizability by generating sharp images with good dynamic range, without any knowledge of the corresponding blur kernel, has no strict limitations on the blur kernel estimation, and is not restrictive to synthetic kernels in our training dataset.
4.4 NSD-DIL for Image Super-Resolution
@@
Bicubic EDSR ZSSR
KernelGAN
+ ZSSR
DBPI DualSR
NSD-LCNN
NSD-DRK
@@
N.A. 2.98 0.29 0.151+0.29 0.5 0.44 0.0028 0.00001
Inference time N.A. 1 =\>== 10 >=>=> =13 >=>=> =1 >=>=> =3 0.002 0.0005
x2 0.7846/27.24/4.127 0.8034/29.17/4.142 0.7925/27.51/4.111 0.8379/28.24/4.071 0.8684/30.77/4.049 0.8250/25.78/5.005 0.8980/27.53/4.976 0.9038/27.27/ 5.006
x4 0.6478/23.89/4.147 0.6615/25.63/4.181 0.6550/24.05/4.156 0.6799/24.76/4.089 0.7368/26.86/4.146 -/-/- 0.7331/23.63/4.904 0.7487/23.93/4.903
@@
Image Super Resolution (ISR) is a well-established low-level vision task whose objective is to generate a High-Resolution (HR) image from the given corresponding LR observation(s). We propose to extend the proposed model to the ISR task. The NSD-DIL model can be directly employed at the inference stage, as the trained L-CNN acts as a pre-upsampled network that operates on traditional upsampled (Bicubic) input blur images of any desired scale factor to generate sharp images with fine details. The same trained L-CNN for the deblurring task was employed to perform the ISR task. We also note that this DRK alone could also be employed for the ISR task in place of L-CNN on a pre-upsampled image and operates the same as L-CNN. In these experiments, no kernel estimation stage is required. The super-resolution ability of our approach was evaluated on the benchmark dataset DIV2KRK [2] for sf 2, and sf 4 using PSNR, SSIM, and NIMA metrics. The ISR results were presented in Table 2. For a fair comparison, we considered works like KernelGAN [2], ZSSR [49], DBPI [20], and DualSR [13] which consider single input LR image only for modeling the ISR task as in our experiments. From the quantitative results, it was observed that the NSD-DIL model delivers competitive performance very close to the SotA ISR works despite being absolutely independent of LR and HR image space. The sample visual results for sf 2 were provided in Fig. 5. The computational complexity in terms of parameters and super-resolution time (inference time) in minutes, of the NSD-DIL model was compared to SotA methods and is given in Table 2. The potential of the proposed NSD-DIL model is observed from the given Table 2. Here, the NSD-DIL requires approximately 100 times fewer memory resources even when compared with the least among the SotA methods i.e. ZSSR, where the DRK requires by order 10 times further lesser than NSD-DIL. Besides, the inference time required for the NSD-DIL and the DRK is reduced by order of 10 and 100 respectively compared to existent zero-shot ISR methods.
@@ 4.5 Ablation study
In this section, the significance of the proposed regularization term ($R$) (refer Eq. 5), and the deblurring performance of the NSD-DIL method on various blur levels along with varied kernel sizes and various sizes of the RKG dataset were discussed. Further, we present the comparison of the deblurring performance of our proposed method with the traditional deconvolution filter i.e. Wiener filter.
@@ 4.5.1 Regularization term (R)
Table 3: The effect of proposed regularization terms in R on the NSD-DIL model’s performance with the synthetic Gaussian blur dataset.
Loss (L) $\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ +
$\lambda_{1}\times R_{1}$
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ +
$\lambda_{2}\times R_{2}$
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ +
$\lambda_{3}\times R_{3}$
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ +
$\lambda_{2}\times R_{2}+\lambda_{3}\times R_{3}$
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ +
$\lambda_{1}\times R_{1}+\lambda_{2}\times R_{2}$
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ +
$\lambda_{1}\times R_{1}+\lambda_{2}\times R_{2}+\lambda_{3}\times R_{3}$
@@
[…]
Note: The deblurring performance of the NSD-DIL method on various blur levels, which are disjoint with our training dataset, was studied and observed only an approximately $6\%$ decrease in terms of PSNR and SSIM metrics as the blur level varied from $\sigma_{1},\sigma_{2}\in U[0.175,3]$ to $U[6,9]$. More details on the study of the deblurring performance of the NSD-DIL method on various blur levels along with varied kernel sizes and various sizes of the RKG dataset details are provided in the supplementary material. In addition, we observed that the conventional inverse degradation modeling by the Wiener filter failed to hold the robustness across the wide range of degradations in the generated simulated blur dataset, itself. The quantitative results are provided in the supplementary material.
4.6 Discussion
In this work, it is important to note that our null-shot model detours the degradation kernel estimation stage to achieve the deblur and ISR tasks, unlike the existing well-established methods including the zero-shot self-supervised deblur method i.e. Polyblur [10] and ISR method i.e. KernelGAN [2].
Here, although we didn’t consider the additive noise explicitly in our degradation model represented by Eq. 1, the proposed model showcased its ability to generate super-resolved images with superior SSIM scores (refer to Table 1 and Table 2). However, in future work, we would extend our null-shot method to address stronger and non-uniform blurs along with additive noise.
In the case of the ISR task, for higher scale factors, the proposed method can deliver improved performance as the inverse degradation kernel size is made large enough. It is planned to explore in our future work on the null-shot approach for other image restoration tasks.
4.7 Conclusion

### v2

@@ 4.1 Training setup
We trained the proposed L-CNN (refer Sec.3.3 for architecture details of L-CNN) on the RKG dataset (refer Sec. 3.2) with the learning objective given in Eq. (4). The number of epochs was $40$ and the learning rate was $10e^{-4}$. The Adam [60] optimization was used, with $\beta=0.9$. The values of hyper-parameters used in Eq. (5), set empirically, are as follows, $\lambda_{1}=0.8,~{}\lambda_{2}=0.8,~{}\lambda_{3}=0.4$. The $D^{3}$ code and RKG dataset will be made publicly available soon for reproducibility. We have used a computer with an NVIDIA-GTX 2080 Ti 11GB GPU in all our experiments.
@@ 4.2 Computational complexity:
The proposed L-CNN is an efficient and lightweight deblurring model. The computational complexity in terms of parameters (for DL based approaches) and inference time of the $D^{3}$ model was compared to SotA methods and is given in Table 1. Here, the proposed $D^{3}$ model requires very less computational resources, by order of 1000 in terms of parameters and inference time when compared to the recent SotA work i.e. IFFS [61]. Besides $D^{3}$, the deblurring task with DRK further reduces the required computational resources parameters and inference time by an order of 10 to $D^{3}$ and by an order of 10K to SotA methods. Thus the proposed $D^{3}$ model is more suitable for real-time deblurring tasks.
@@ 4.3 Results
The deblurring performance of our proposed method is compared with a wide range of methods including standard and recent non-DL blind deblurring methods NSM [5], 1Shot-MaxPol [38], Guided Image Filtering (GIF) [62], PIBKE [63], Polyblur [1] and DL-based methods DeblurGAN [27], USRNet [26], XYDeblur [64], IFFS [61]. We note that the proposed $D^{3}$ is only the DL approach among the above-mentioned DL-based methods that achieve the deblurring task without training on image data.
@@
Method Approach
@@
(in seconds)
PSNR SSIM
NSM Supervised/Non-DL N.A. 120 27.51 0.8170
PIBKE Self-supervised/Non-DL N.A. 120 24.86 0.7384
GIF Supervised/Non-DL N.A. 60 19.09 0.6614
1Shot-MaxPol Unsupervised/Non-DL N.A. 90 25.88 0.7851
Deblur GAN Supervised/DL 17.1 1 26.65 0.7933
USRNet Unsupervised/DL 6.1 1 24.31 0.7396
XYDeblur Supervised/DL 4.92 0.729 27.18 0.8059
Polyblur Self-supervised/Non-DL N.A. 0.001 27.10 0.7794
IFFS Supervised/DL 19.5 3.5 26.52 0.8242
NSSR-DIL
Image-data
independent/DL
0.0028 0.002 25.19 0.8133
$D^{3}$
Image-data
independent/DL
0.0028 0.002 27.01 0.8333
$D^{3}$
Image-data
independent/DL
0.00001 0.0005 28.02 0.8383
@@
Deblur performance comparison: We generated a synthetic blur dataset following the [1]. Random Gaussian kernels of varying shapes, orientations (i.e. $\sigma_{1},\sigma_{2}\in U[0.175,3]$ and rotation angle $\theta\in U[0,\pi]$) and sizes $11\times 11,15\times 15,19\times 19,23\times 23,27\times 27$ were generated. Each set of kernels, varying in size, was applied to the set of 100 sharp images from the DIV2K validation dataset [65], resulting in 500 test images. To have a little deviation from the Gaussian distribution, a small amount of multiplicative noise was introduced to the random Gaussian kernels. We compared the $D^{3}$ model to the above-mentioned SotA works using the simulated blur dataset. To assess the ability to restore the finer details during the deblurring process, we employ PSNR and SSIM for qualitative and quantitative similarity with the corresponding sharp images. The results were tabulated and presented in Table 1. The qualitative results were provided in Fig. 4.
Despite being very lightweight, the proposed $D^{3}$ demonstrated its effectiveness significantly with standard metrics. The compared SotA methods need supervised learning from images for $K$ estimation and/or deblurring model learning. Thus requires more memory, and time resources at least by order of 100 in parameters and inference time respectively compared to the proposed $D^{3}$.
Evaluation on real data: Further, we provide the qualitative comparison of deblurring performance on real captured blurry images provided in Fig. 1 and Fig.5. The qualitative results given in Fig. 4, Fig. 1 and Fig.5 illustrate that the proposed model did not produce artifacts in the generated deblurred output images, unlike the compeer, zero-shot non-DL SotA works like [5], [63]. Besides, the proposed $D^{3}$ method demonstrates its generalizability by generating sharp images with good dynamic range, without any knowledge of the corresponding blur kernel, has no strict limitations on the blur kernel estimation, and is not restrictive to synthetic kernels in our training dataset.
Figure 5: Deblurring performance of various methods on real captured images using smartphone Samsung J7 Prime (13MP, f/1.9)
4.4 $D^{3}$ for Image Super-Resolution
@@
Method
Scale
factor
Inference
time
(in seconds)
@@
(in Millions)
SSIM PSNR NIMA
Bicubic X2 N.A. N.A. 0.7846 27.24 4.127
X4 0.6478 23.89 4.147
ZSSR X2 0.29 =10 0.7925 27.51 4.111
X4 0.6550 24.05 4.156
KernelGAN +ZSSR X2 0.151+0.29 >=13 0.8379 28.24 4.071
X4 0.6799 24.76 4.089
DBPI X2 0.5 >=1 0.8684 30.77 4.049
X4 0.7368 26.86 4.146
DualSR X2 0.44 >=3 0.8250 25.78 5.005
X4 - - -
NSSR-DIL X2 0.0028 0.002 0.8644 26.02 4.161
X4 0.7926 23.58 4.170
$D^{3}$ (LCNN) X2 0.0028 0.002 0.8980 27.53 4.976
X4 0.7331 23.63 4.904
$D^{3}$ (DRK) X2 0.00001 0.0005 0.9038 27.27 5.006
X4 0.7487 23.93 4.903
@@
Image Super Resolution (ISR) is a well-established low-level vision task whose objective is to generate a High-Resolution (HR) image from the given corresponding LR observation(s). We propose to extend the proposed model to the ISR task. The $D^{3}$ model can be directly employed at the inference stage, as the trained L-CNN acts as a pre-upsampled network that operates on traditional upsampled (Bicubic) input blur images of any desired scale factor to generate sharp images with fine details. The same trained L-CNN for the deblurring task was employed to perform the ISR task. We also note that this DRK alone could also be employed for the ISR task in place of L-CNN on a pre-upsampled image and operates the same as L-CNN. In these experiments, no kernel estimation stage is required. The super-resolution ability of our approach was evaluated on the benchmark dataset DIV2KRK [66] for sf 2, and sf 4 using PSNR, SSIM, and NIMA metrics. The ISR results were presented in Table 2. For a fair comparison, we considered works like KernelGAN [66], ZSSR [67], DBPI [68], and DualSR [69] which consider single input LR image only for modeling the ISR task as in our experiments. From the quantitative results, it was observed that the $D^{3}$ model delivers competitive performance very close to the SotA ISR works despite being absolutely independent of LR and HR image space. The sample visual results for sf 2 were provided in Fig. 6. The computational complexity in terms of parameters and super-resolution time (inference time) in minutes, of the $D^{3}$ model was compared to SotA methods and is given in Table 2. The potential of the proposed $D^{3}$ model is observed from the given Table 2. Here, the $D^{3}$ requires approximately 100 times fewer memory resources even when compared with the least among the SotA methods i.e. ZSSR, where the DRK requires by order 10 times further less than $D^{3}$. Besides, the inference time required for the $D^{3}$ and the DRK is reduced by order of 10 and 100 respectively compared to existent zero-shot ISR methods.
Notably, while both our current proposed method and previous work (i.e. NSSR-DIL) utilize the L-CNN model, the $D^{3}$ framework exhibits marked gains in deblurring and ISR performance compared to our earlier approach, primarily attributable to the updated regularization term.
@@ 4.5 Ablation study
In this section, the significance of the proposed regularization term ($R$) (refer Eq. 5), and the deblurring performance of the $D^{3}$ method on various blur levels along with varied kernel sizes and various sizes of the RKG dataset were discussed. Further, we present the comparison of the deblurring performance of our proposed method with the traditional deconvolution filter i.e. Wiener filter.
@@ 4.5.1 Regularization term (R)
Table 3: The effect of proposed regularization terms in R on the $D^{3}$ model’s performance with the synthetic Gaussian blur dataset.
@@
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}$ 0.0224 6.59
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}+\lambda_{1}\times L_{ConvArea}$ 0.5455 21.07
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}+\lambda_{2}\times L_{PhLoss}$ 0.2137 8.12
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}+\lambda_{3}\times L_{MagLoss}$ 0.2030 17.69
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}+\lambda_{2}\times L_{PhLoss}+\lambda_{3}\times L_{MagLoss}$ 0.5644 17.21
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}+\lambda_{1}\times L_{ConvArea}+\lambda_{2}\times L_{PhLoss}$ 0.7419 25.04
$\lVert K\ast K^{-1}-~{}\delta\rVert^{2}_{2}+\lambda_{1}\times L_{ConvArea}+\lambda_{2}\times L_{PhLoss}+\lambda_{3}\times L_{MagLoss}$ 0.8333 27.01
4.5.2 Evaluation on various blur levels:
We evaluated the deblurring performance of the proposed $D^{3}$ model on various blur levels, whose blur kernels are disjoint with our training dataset. We synthetically generated a blur dataset following similar techniques used to generate a blur test dataset as discussed in Sec. 4 with different sets of variances $\sigma_{1},\sigma_{2}\in U[0.175,3],U[3,6],~{}\&~{}U[6,9]$ and using different kernel sizes i.e. $11\times 11,15\times 15,19\times 19~{}and~{}21\times 21$. The results were outlined in Table 4. $D^{3}$ model trained on $11\times 1$ 11 size kernels, demonstrates robustness even when tested against a broad spectrum of blurs, encompassing varied kernel sizes such as $11\times 11,15\times 15,19\times 19~{}and~{}21\times 21$. The experimental findings showcase the overall effectiveness and satisfactory deblurring capabilities of the $D^{3}$ model.
Table 4: Deblurring performance of the proposed $D^{3}$ model on various blur levels using PSNR/SSIM metrics.
Degradation kernel $K$ size
$11\times 1$ 11 $15\times 15$ $19\times 19$ $21\times 21$
Variance levels $\sigma_{1},\sigma_{2}\in U[0.1,3]$ 26.67/0.8436 26.53/0.8451 26.53/0.8441 26.61/0.8440
$\sigma_{1},\sigma_{2}\in U[3,6]$ 25.92/0.7726 25.89/0.7727 25.79/0.7697 25.84/0.7720
$\sigma_{1},\sigma_{2}\in U[6,9]$ 25.02/0.7220 24.74/0.7119 24.72/0.7124 24.68/0.7102
4.5.3 Performance of $D^{3}$ with varied sizes of RKG dataset
We studied the performance of the $D^{3}$ model trained on varied sizes of the RKG dataset to empirically find the size of the RKG dataset to train the proposed $D^{3}$ model. In our study, we generated five different datasets with the number of samples $800,1600,2400,3200,4000$, and $4800$ and named as $RKG_{n}$, where $n=1,2,3,4,5$, respectively. These five datasets i.e. $RKG_{n}$ were generated for kernel sizes $11\times 11~{}\&~{}21\times 21$, following the steps discussed in the case of RKG dataset generation. The L-CNN was trained independently on each set and evaluated on the synthetic Gaussian blur dataset using PSNR and SSIM metrics. The visual plots depicting the L-CNN performance comparison for $RKG_{n}$ vs evaluation metrics i.e. PSNR, and SSIM are presented in Fig. 7 (a) and Fig. 7 (b) respectively. It is observed that there is a linear improvement in these metrics as the number of samples was increased initially and later fell off the clip. Therefore we considered 2400 training samples i.e. $RKG_{3}$ dataset, with kernel size $11\times 1$ 11, in all our experiments.
Figure 7: The performance comparison of the $D^{3}$ model with sizes of RKG dataset.
4.5.4 Deblurring performance comparison of the proposed method with Wiener filter.
In our experiments, we compared the deblurring performance of our proposed learnable deconvolution kernel with the traditional inverse filter, i.e. the Wiener filter. We observed that the performance of the Wiener filter is not robust across the dataset and shows $>$50% poor values in evaluation metrics, PSNR, and SSIM. We also trained a Single Layer and Single Channel (SLSC) linear CNN, to compute the generalized inverse kernel for the RKG dataset using the DIL objective and the proposed regularization term. The results were tabulated in Table 5. In contrast to the performance of traditional Wiener filter’s and SLSC networks, the proposed method can handle a wide range of degradations across the dataset and generate deblurred images with fine details.
Table 5: Comparison of deblurring performance of $D^{3}$ and Wiener filter on the generated synthetic blur dataset, discussed in Section 4.3.
Method PSNR SSIM
Wiener Filter 10.49 0.1752
$D^{3}$ (L-CNN-SLSC) 19.28 0.5248
$D^{3}$ (L-CNN) 27.01 0.8333
$D^{3}$ (DRK) 28.02 0.8383
5 Conclusion